40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 1 15:46:12 2017, Kira, Update, PEM, temp sensor update IMG_20170901_144729.jpg
    Reply  Tue Sep 5 17:18:17 2017, Kira, Update, PEM, temp sensor update IMG_20170905_144924.jpgIMG_20170905_165042.jpg
       Reply  Tue Sep 5 17:52:06 2017, Kira, Update, PEM, temp sensor update 
          Reply  Fri Sep 8 10:22:30 2017, Steve, Update, PEM, temp sensor update 
       Reply  Tue Sep 12 11:44:16 2017, Kira, Update, PEM, temp sensor update 
Message ID: 13285     Entry time: Fri Sep 1 15:46:12 2017     Reply to this: 13295
Author: Kira 
Type: Update 
Category: PEM 
Subject: temp sensor update 

I took off the AD590 and attached it to two long wires leading out from the board. This will allow us to attach the sensor to a metal block and not have to stick the whole board to it. I have also completed three identical copies of this and it's pretty much ready to be tested. According to Craig and Andrew's elog here, the sensor is very noisy and they added in a low pass filter to fix that, so that's something to consider for the final version of the circuit. I'll test what I have so far and see how that goes. We still need to figure out how to get readings from the sensors.

To attach the sensor to the metal block, I'll use some thermal paste and fasteners. I'll also put a thermometer on the block to record the actual temperature. I'll then wrap it in some insulation we have in the lab and have only some wires leading out of it to make measurements. I'll leave this setup overnight and record the outputs for about a full day. The fluctuations between the sensors will then indicate the noise of each individual sensor.

Attachment 1: IMG_20170901_144729.jpg  2.670 MB  | Hide | Hide all
IMG_20170901_144729.jpg
ELOG V3.1.3-